# Twisted character formula for toral supercuspidal representations

### v1

@@ 6.6. Twisted characters of Weil representations: summary
In all cases, $\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})$ is given by
$\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})=\epsilon_{\Sigma\alpha}([\eta]^{m_{\alpha}})\cdot|V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}},$
where $\epsilon_{\Sigma\alpha}([\eta]^{m_{\alpha}})$ is a sign depending on the root $\alpha$ and the element $\eta]^{m_{\alpha}}$ as described in Sections 6.4 and 6.5. We did not explicate the sign $\epsilon_{\Sigma\alpha}([\eta]^{m_{\alpha}})$ in general, so here let us state the result just in the following way:
@@
$\prod_{\alpha\in\Theta_{\mathbf{S}}\backslash\ddot{\Xi}}\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})=|V_{\eta}|^{\frac{1}{2}}\cdot\prod_{\alpha\in\Theta_{\mathbf{S}}\backslash\ddot{\Xi}}\epsilon_{\Sigma\alpha}([\eta]^{m_{\alpha}}).$
However, recall that our computation of each sign factor $\epsilon_{\Sigma\alpha}([\eta]^{m_{\alpha}})$ becomes quite simple in the case where $f_{\alpha}=1$. From a practical point of view (especially, towards an application to the twisted endoscopic character relation), let us rewrite the formula of Proposition 6.12 in this case where $f_{\alpha}=1$ holds for all $\alpha$.
We introduce characters $\epsilon_{\alpha}\colon S\rightarrow\mathbb{C}^{\times}$ for $\alpha\in\Phi(\mathbf{G},\mathbf{S})$ as follows:
@@
where $\epsilon_{\Sigma\alpha}$ denotes the sign only determined by $\alpha$ in the case where $\alpha_{\mathrm{res}}$ is symmetric ramified.
@@
$\epsilon_{\Sigma\alpha}([\eta]^{m_{\alpha}})=\operatorname{sgn}_{\mathbb{F}_{p}^{\times}}(-1)^{g_{\alpha}(f_{\alpha}-1)}\cdot\operatorname{sgn}_{k_{\alpha}^{\times}}(\underline{\eta}_{\alpha})\cdot\prod_{i=0}^{m_{\alpha}-1}\epsilon_{\theta^{i}(\alpha)}(s).$
@@
$\epsilon_{\Sigma\alpha}([\eta]^{m_{\alpha}})=(-1)^{n_{\alpha}(\eta)}\cdot\operatorname{sgn}_{k_{\alpha}^{\times}}(\underline{\eta}_{-\alpha}C)\cdot\prod_{i=0}^{m_{\alpha}-1}\epsilon_{\theta^{i}(\alpha)}(s).$
@@
$\epsilon_{\Sigma\alpha}([\eta]^{m_{\alpha}})=(-1)^{1-n_{\alpha}(\eta)}\cdot\operatorname{sgn}_{k_{\alpha}^{1}}(\underline{\eta}_{\alpha})\cdot\prod_{i=0}^{m_{\alpha}-1}\epsilon_{\theta^{i}(\alpha)}(s).$
We define a constant $\epsilon_{\Sigma\alpha}$ for $\alpha$ whose $\alpha_{\mathrm{res}}$ is asymmetric or symmetric ramified by
@@
(Note that this constant depends only on $\alpha$ and the choice of the fixed base point $\underline{\eta}$ and does not depend on $\eta$.) We define a constant $C_{\underline{\eta}}$ to be the product of $\epsilon_{\Sigma\alpha}$’s:
$C_{\underline{\eta}}:=\prod_{\alpha\in\Theta_{\mathbf{S}}\backslash\ddot{\Xi}}\epsilon_{\Sigma\alpha}.$
@@
$\tilde{\epsilon}_{\Xi}(s)\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\prod_{\begin{subarray}{c}\alpha\in\ddot{\Xi}\\ \alpha_{\mathrm{res}}:\,\mathrm{asym}/\mathrm{ur}\end{subarray}}\epsilon_{\alpha}(s).$
@@
$\prod_{\alpha\in\Theta_{\mathbf{S}}\backslash\ddot{\Xi}}\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})=C_{\underline{\eta}}\cdot(-1)^{|\ddot{\Xi}_{\eta,\mathrm{ur}}|}\cdot|V_{\eta}|^{\frac{1}{2}}\cdot\tilde{\epsilon}_{\Xi}(s).$
@@
$\Theta_{\tilde{\rho}}(\eta)=|V_{\eta_{0}}|^{\frac{1}{2}}\cdot\prod_{\alpha\in\Theta_{\mathbf{S}}\backslash\ddot{\Xi}}\Theta_{\omega_{\Sigma\alpha}}([\eta_{0}]^{m_{\alpha}})\cdot\vartheta(s).$
@@
$\operatorname{tr}\bigl(\omega([s])\circ I_{\omega}^{\underline{\eta}}\bigr)=|V_{\eta_{0}}|^{\frac{1}{2}}\cdot\prod_{\alpha\in\Theta_{\mathbf{S}}\backslash\ddot{\Xi}}\Theta_{\omega_{\Sigma\alpha}}([\eta_{0}]^{m_{\alpha}}).$
@@
$\prod_{\alpha\in\Theta_{\mathbf{S}}\backslash\ddot{\Xi}}\Theta_{\omega_{\Sigma\alpha}}([\eta_{0}]^{m_{\alpha}})=C_{\underline{\eta}}\cdot(-1)^{|\ddot{\Xi}_{\eta_{0},\mathrm{ur}}|}\cdot\tilde{\epsilon}_{\Xi}(s^{\prime})$
@@ 7. Twisted Adler–DeBacker–Spice character formula
$\Theta_{\tilde{\rho}}({}^{g}\eta)=|V_{{}^{g}\eta_{0}}|^{\frac{1}{2}}\cdot\prod_{\alpha\in\Theta_{\mathbf{S}}\backslash\ddot{\Xi}}\Theta_{\omega_{\Sigma\alpha}}([{}^{g}\eta_{0}]^{m_{\alpha}})\cdot\vartheta(s_{g}).$

### v2

@@ 6.6. Twisted characters of Weil representations: summary
In all cases, $\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})$ is given by
$\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})=\epsilon_{\ddot{\alpha}}([\eta]^{m_{\alpha}})\cdot|V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}},$
where we let $\epsilon_{\ddot{\alpha}}([\eta]^{m_{\alpha}})$ denote the sign depending on the root $\alpha$ and the element $\eta]^{m_{\alpha}}$ as described in Sections 6.4 and 6.5. We did not explicate the sign $\epsilon_{\ddot{\alpha}}([\eta]^{m_{\alpha}})$ in general, so here let us state the result just in the following way:
@@
$\prod_{\ddot{\alpha}\in\Theta_{\mathbf{S}}\backslash\ddot{\Xi}}\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})=|V_{\eta}|^{\frac{1}{2}}\cdot\prod_{\ddot{\alpha}\in\Theta_{\mathbf{S}}\backslash\ddot{\Xi}}\epsilon_{\ddot{\alpha}}([\eta]^{m_{\alpha}}).$
However, recall that our computation of each sign factor $\epsilon_{\ddot{\alpha}}([\eta]^{m_{\alpha}})$ becomes quite simple in the case where $f_{\alpha}=1$. From a practical point of view (especially, towards an application to the twisted endoscopic character relation), let us rewrite the formula of Proposition 6.12 in this case where $f_{\alpha}=1$ holds for all $\alpha$.
We introduce characters $\epsilon_{\alpha}\colon S\rightarrow\mathbb{C}^{\times}$ for $\alpha\in\Phi(\mathbf{G},\mathbf{S})\smallsetminus\Phi(\mathbf{G},\mathbf{S})_{\mathrm{ram}}$ as follows:
@@
where $\epsilon_{\ddot{\alpha}}$ denotes the sign only determined by $\alpha$ in the case where $\alpha_{\mathrm{res}}$ is symmetric ramified.
@@
$\epsilon_{\ddot{\alpha}}([\eta]^{m_{\alpha}})=\operatorname{sgn}_{\mathbb{F}_{p}^{\times}}(-1)^{g_{\alpha}(f_{\alpha}-1)}\cdot\operatorname{sgn}_{k_{\alpha}^{\times}}(\underline{\eta}_{\alpha})\cdot\prod_{i=0}^{m_{\alpha}-1}\epsilon_{\theta^{i}(\alpha)}(s).$
@@
$\epsilon_{\ddot{\alpha}}([\eta]^{m_{\alpha}})=(-1)^{n_{\alpha}(\eta)}\cdot\operatorname{sgn}_{k_{\alpha}^{\times}}(\underline{\eta}_{-\alpha}C)\cdot\prod_{i=0}^{m_{\alpha}-1}\epsilon_{\theta^{i}(\alpha)}(s).$
@@
$\epsilon_{\ddot{\alpha}}([\eta]^{m_{\alpha}})=(-1)^{1-n_{\alpha}(\eta)}\cdot\operatorname{sgn}_{k_{\alpha}^{1}}(\underline{\eta}_{\alpha})\cdot\prod_{i=0}^{m_{\alpha}-1}\epsilon_{\theta^{i}(\alpha)}(s).$
We define a constant $\epsilon_{\ddot{\alpha}}$ for $\alpha$ whose $\alpha_{\mathrm{res}}$ is asymmetric or symmetric ramified by
@@
(Note that this constant depends only on $\alpha$ and the choice of the fixed base point $\underline{\eta}$ and does not depend on $\eta$.) We define a constant $C_{\underline{\eta}}$ to be the product of $\epsilon_{\ddot{\alpha}}$’s:
$C_{\underline{\eta}}:=\prod_{\ddot{\alpha}\in\Theta_{\mathbf{S}}\backslash\ddot{\Xi}}\epsilon_{\ddot{\alpha}}.$
@@
$\tilde{\epsilon}_{\Xi}(s)\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\prod_{\begin{subarray}{c}\ddot{\alpha}\in\ddot{\Xi}\\ \alpha_{\mathrm{res}}:\,\mathrm{asym}/\mathrm{ur}\end{subarray}}\epsilon_{\alpha}(s).$
@@
$\prod_{\ddot{\alpha}\in\Theta_{\mathbf{S}}\backslash\ddot{\Xi}}\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})=C_{\underline{\eta}}\cdot(-1)^{|\ddot{\Xi}_{\eta,\mathrm{ur}}|}\cdot|V_{\eta}|^{\frac{1}{2}}\cdot\tilde{\epsilon}_{\Xi}(s).$
@@
$\Theta_{\tilde{\rho}}(\eta)=|V_{\eta_{0}}|^{\frac{1}{2}}\cdot\prod_{\ddot{\alpha}\in\Theta_{\mathbf{S}}\backslash\ddot{\Xi}}\Theta_{\omega_{\ddot{\alpha}}}([\eta_{0}]^{m_{\alpha}})\cdot\vartheta(s).$
@@
$\operatorname{tr}\bigl(\omega([s])\circ I_{\omega}^{\underline{\eta}}\bigr)=|V_{\eta_{0}}|^{\frac{1}{2}}\cdot\prod_{\ddot{\alpha}\in\Theta_{\mathbf{S}}\backslash\ddot{\Xi}}\Theta_{\omega_{\ddot{\alpha}}}([\eta_{0}]^{m_{\ddot{\alpha}}}).$
@@
$\prod_{\ddot{\alpha}\in\Theta_{\mathbf{S}}\backslash\ddot{\Xi}}\Theta_{\omega_{\ddot{\alpha}}}([\eta_{0}]^{m_{\ddot{\alpha}}})=C_{\underline{\eta}}\cdot(-1)^{|\ddot{\Xi}_{\eta_{0},\mathrm{ur}}|}\cdot\tilde{\epsilon}_{\Xi}(s^{\prime})$
@@ 7. Twisted Adler–DeBacker–Spice character formula
$\Theta_{\tilde{\rho}}({}^{g}\eta)=|V_{{}^{g}\eta_{0}}|^{\frac{1}{2}}\cdot\prod_{\ddot{\alpha}\in\Theta_{\mathbf{S}}\backslash\ddot{\Xi}}\Theta_{\omega_{\ddot{\alpha}}}([{}^{g}\eta_{0}]^{m_{\ddot{\alpha}}})\cdot\vartheta(s_{g}).$